OpenStudy (anonymous):

integral 1+x^2/1+x^4 from 0 to infinity

OpenStudy (lgbasallote):

\[\Large \int \limits_0^\infty \frac{1+x^2}{1+x^4}dx\] is that the question?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

let\[t=x-\frac{1}{x}\]

OpenStudy (anonymous):

\[\int_{0}^{\infty}\frac{1+x^2}{1+x^4} \text{d}x=\int_{-\infty}^{\infty}\frac{1}{t^2+2} \text{d}t\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

oh yes...thank u

OpenStudy (anonymous):

yw :)